NauticalNautical

​& Co. owns vast amounts of corporate bonds. Suppose

NauticalNautical

buys

$ 600 comma 000$600,000

of

BitterCoBitterCo

bonds at face value on January​ 2,

20182018.

The

BitterCoBitterCo

bonds pay interest at the annual rate of

77​%

on June 30 and December 31 and mature on December​ 31,

20322032.

NauticalNautical

intends to hold the investment until maturity.

Requirements

1.

Journalize any required

20182018

entries for the bond investment.

2.

How much cash interest will

NauticalNautical

receive each year from

BitterCoBitterCo​?

3.

How much interest revenue will

NauticalNautical

report during

20182018

on this bond​ investment?

Requirement 1. Journalize any required

20182018

entries for the bond investment. ​(Record debits​ first, then credits. Select the explanation on the last line of the journal entry​ table.)Begin by journalizing

NauticalNautical​'s

investment on January​ 2,

20182018.

Date

Accounts and Explanation

Debit

Credit

2018

Jan. 2

​Next, journalize the receipt of cash interest on June​ 30,

20182018.

Date

Accounts and Explanation

Debit

Credit

2018

Jun. 30

                 Now journalize the receipt of cash interest on December​ 31,

20182018.

Date

Accounts and Explanation

Debit

Credit

2018

Dec. 31

Requirement 2. How much cash interest will

NauticalNautical

receive each year from

BitterCoBitterCo​?

Nautical will receive $

of cash interest each year from BitterCo.

Requirement 3. How much interest revenue will

NauticalNautical

report during

20182018

on this bond​ investment?

Nautical will report $

of interest revenue during 2018 on this bond investment.

Choose from any list or enter any number in the input fields and then continue to the next question.
Answer #1
1 Date Accounts and explanation Debit Credit
2018
Jan 2. Investment in Bitter co bonds 600000
Cash 600000
(Invested in bonds)
June 30. Cash (600000*7%*6/12) 21000
Interest revenue 21000
(Interest received on bonds)
Dec 31. Cash (600000*7%*6/12) 21000
Interest revenue 21000
(Interest received on bonds)
2 Cash interest per year=600000*7%=$ 42000
3 Interest revenue for 2018=21000+21000=$ 42000